# Stock Market Prediction Using Node Transformer Architecture Integrated with BERT Sentiment Analysis

### 06 Mar 2026 v1

@@ V-E4 Performance Across Volatility Regimes
The proposed model maintains MAPE below 1.5% during high-volatility periods while baseline models exceed 1.6-2.1%. This robustness gap is most pronounced during VIX spikes accompanying earnings surprises and macroeconomic shocks, precisely the conditions where accurate forecasts carry the greatest practical value for risk management. The gating mechanism likely contributes to this resilience by increasing the weight of momentum features during volatile periods while favoring longer-term indicators during calm markets.
V-E5 Statistical Significance
@@
V-E6 Economic Significance
A long-short strategy going long on the top 5 predicted performers and short on the bottom 5 achieved gross cumulative returns of 25.3% over the 15-month evaluation period (January 2024 to March 2025), compared to 15.1% for S&P 500 buy-and-hold.
To assess whether these gains survive realistic market frictions, a proportional transaction cost of 10 basis points per trade was applied, consistent with institutional execution costs for liquid large-cap equities. The strategy exhibited an average daily portfolio turnover of 32%, reflecting the sentiment-driven rebalancing that shifts positions as new social media signals arrive. After deducting transaction costs, the net cumulative return was 18.4%, still exceeding the buy-and-hold benchmark by 3.3 percentage points. The net annualized Sharpe ratio of 1.15 (down from 1.42 gross) confirms that the model’s directional accuracy translates into viable risk-adjusted gains even under trading friction, though the sensitivity to turnover suggests that lower-frequency rebalancing schedules merit investigation for cost-sensitive applications.
@@ VI-A Interpretation of Results
The integrated model demonstrates consistent improvements across metrics, horizons, and market conditions, with the magnitude of gains varying by component and context in ways that illuminate the mechanisms underlying the model’s predictive advantage.
@@
Sentiment integration proves valuable around information events. The 25% improvement during earnings seasons indicates that social media sentiment contains early signals not fully reflected in contemporaneous prices. Sector-level heterogeneity aligns with expectations: consumer-facing companies show stronger sentiment-price relationships (14% MAPE reduction) compared to utilities (6%), likely reflecting differential retail investor engagement.
@@ VI-B Implications
For practitioners, the model’s directional accuracy enables more informed trading decisions, particularly for short-term strategies where predicting price direction is often more important than predicting magnitude. The enhanced ability to predict during volatile conditions helps risk management, as these are precisely the periods when forecast reliability matters most for position sizing and drawdown control. The graph structure provides interpretable inter-stock relationships useful for portfolio diversification, as the learned edge weights can reveal dependencies that traditional correlation matrices might obscure.
From a theoretical perspective, performance gains from sentiment integration suggest that markets do not fully incorporate qualitative information immediately, supporting bounded efficiency [1]. The 25% improvement during earnings seasons indicates that social media contains early signals about investor reactions that are not yet reflected in price. The success of graph-based modeling provides further evidence that markets function as complex, interconnected systems rather than collections of independent securities, consistent with findings in the network finance literature.
@@ VI-C Limitations
The most consequential concern is survivorship bias in stock selection. The 20 stocks were chosen based on current S&P 500 membership and data availability, meaning that companies which failed, were acquired, or were delisted between 1982 and 2025 are absent from the dataset. Because the selected universe consists of companies that ultimately succeeded, reported performance may be inflated relative to what a real-time investor would experience. Results could differ on a point-in-time universe constructed from historical index constituents. Relatedly, the 20-node graph structure, while demonstrating meaningful contribution in ablation studies (15% improvement), represents a limited cross-section of the market. Larger universes might better capture complex market topology; the current design prioritizes temporal depth over cross-sectional breadth.
Beyond stock selection, data-related assumptions also affect generalizability. Edge weight initialization relies on correlations computed over the training period (1982-2010), but market correlations are inherently non-stationary. Relationships that held during this window may weaken or reverse by the test period. The learnable edge refinement mechanism partially compensates by adapting weights during training, though fundamental correlation regime shifts could still limit generalization. Similarly, X (formerly Twitter) data availability from 2007 onward means sentiment features are set to zero for the 1982-2006 training period. The sentiment component therefore receives limited training signal during most of the training window, with the model’s ability to leverage sentiment primarily developed and evaluated on post-2007 data. Sentiment analysis also relies on English-language content from a single social media platform, and changes to API access policies may affect reproducibility.
@@ VI-D Future Directions
Multiple directions merit investigation. Expanding the stock universe to the full S&P 500 or international markets would test whether the graph-based architecture scales effectively and whether inter-stock dependencies remain informative as the number of nodes grows. Incorporating multilingual sentiment sources and alternative platforms (Reddit, financial news, earnings call transcripts) would broaden the information base and reduce single-source dependency. On the architectural side, developing more efficient attention mechanisms, such as sparse or linear attention variants, could enable real-time deployment for higher-frequency trading applications. Regime-shift detection mechanisms that adjust model behavior during structural market transitions (e.g., policy changes or liquidity crises) represent a natural extension of the gating mechanism already present in the model. Finally, integrating the forecasting framework with portfolio optimization and formal risk management objectives could bridge the gap between prediction accuracy and realized economic value.
@@ VII Conclusion
Taken together, these findings support the premise that jointly modeling inter-stock dependencies and investor sentiment yields more accurate and robust forecasts than architectures that address either dimension in isolation. Future work should extend this framework to broader equity universes, incorporate multilingual and multi-platform sentiment sources, investigate efficient attention variants for real-time deployment, and integrate the forecasting pipeline with portfolio optimization and formal risk management objectives.

### 03 Apr 2026 v2

@@ V-E3 Ablation Studies
The ablation results also address a natural question about the sentiment data horizon. Because social media data begins in 2007 while the financial dataset extends to 1982, the model’s price-based components receive training signal over the full temporal range whereas the sentiment components train on only four years of labeled data (2007-2010). The “Without Sentiment” configuration in Table VII is functionally equivalent to a price-only evaluation of the architecture, achieving MAPE of 0.88% and demonstrating that the node transformer and graph structure are fundamentally competent without any sentiment signal. The full model’s improvement to 0.80% then isolates the incremental contribution of sentiment integration. Because the entire test set (2017-2025) falls within the sentiment-available period, these two configurations effectively provide the decomposition of a long-horizon price-based evaluation alongside a sentiment-augmented evaluation on the same test window, confirming that the 10% reduction represents genuine predictive value rather than an artifact of the training mismatch. Given that the sentiment fusion mechanism was trained on only four years of multimodal data (2007-2010), this improvement likely represents a conservative estimate of the contribution sentiment could offer with a longer training history.
@@ V-E4 Performance Across Volatility Regimes
The proposed model achieves MAPE of 1.50% during high-volatility periods, compared to 1.60–2.10% for baseline models. This robustness gap is most pronounced during VIX spikes accompanying earnings surprises and macroeconomic shocks, precisely the conditions where accurate forecasts carry the greatest practical value for risk management. The gating mechanism likely contributes to this resilience by increasing the weight of momentum features during volatile periods while favoring longer-term indicators during calm markets.
Directional accuracy also varies with market regime. During low-volatility periods (VIX <15<15), the proposed model achieves 68.2% DA, benefiting from the relatively stable trend structure that characterizes calm markets. In medium-volatility conditions ($15\leq$ VIX <25<25), DA decreases to 64.5%, and during high-volatility episodes (VIX $\geq 25$) it falls to 59.8%. While the high-volatility DA represents a meaningful decline from calm-market performance, it remains well above the 50% random baseline and exceeds all baseline models by 3-5 percentage points in the same regime. This pattern is consistent with the expectation that directional prediction becomes inherently more difficult as volatility increases and price movements are dominated by sudden information shocks rather than gradual trends.
V-E5 Return-Based Evaluation
To verify that the reported improvements reflect genuine return prediction rather than an artifact of price persistence, prediction accuracy was also assessed in return space. For each model, the implied daily return is computed as $\hat{r}_{t}=(\hat{y}_{t+1}-y_{t})/y_{t}$ and compared against the realized return $r_{t}=(y_{t+1}-y_{t})/y_{t}$. Because daily price changes for mega-cap equities are small relative to the price level ($|r_{t}|\ll 1$), the return prediction MAE is mathematically close to the price-level MAPE: specifically, return MAE $\text{MAPE}\times(y_{t+h}/y_{t})$, which approaches MAPE as $h\rightarrow 0$. For the 1-day horizon, the proposed model achieves a return MAE of 0.81%, the naive random walk yields 1.36% (reflecting the average absolute daily return of the 20-stock universe), and all intermediate models fall between these bounds in the same rank order as Table V. This near-equivalence between price and return metrics at the daily horizon confirms that the MAPE rankings are not inflated by the choice of evaluation target.
The more informative return-based assessments for this dataset are the metrics already reported: Theil’s U (0.59) explicitly benchmarks against persistence forecasting, directional accuracy (65%) evaluates trend detection independent of magnitude, and the Information Coefficient measures cross-sectional ranking on returns. At longer horizons where the gap between $y_{t}$ and $y_{t+h}$ widens, return-based evaluation becomes more distinct from price-level metrics. The proposed model’s 5-day return MAE of 1.82% and 20-day return MAE of 2.85% remain proportionally below all baselines, with the performance gap widening at longer horizons consistent with the MAPE results.
V-E6 Statistical Significance
@@
V-E7 Economic Significance
To assess practical utility beyond statistical accuracy, a simulated long-short portfolio strategy was constructed over the full test period. Each trading day, the strategy takes long positions in the top 5 predicted performers and short positions in the bottom 5, equally weighted and rebalanced daily. The following results should be interpreted as a preliminary indication of economic viability rather than a definitive backtest, given the simplified assumptions regarding execution and the use of survivor-biased stock selection.
Over the full test period (January 2017 to March 2025), the strategy achieved gross annualized returns of 22.8%, compared to 13.9% for S&P 500 buy-and-hold. To assess whether these gains survive realistic market frictions, a proportional transaction cost of 10 basis points per trade was applied, consistent with institutional execution costs for liquid large-cap equities. The strategy exhibited an average daily portfolio turnover of 32%, reflecting the sentiment-driven rebalancing that shifts positions as new social media signals arrive. After deducting transaction costs, the net annualized return was 14.6%, still exceeding the buy-and-hold benchmark by 0.7 percentage points. The net annualized Sharpe ratio was 0.84 (down from 1.35 gross), and the maximum drawdown reached 14.3%, compared to 23.9% for the buy-and-hold benchmark during the same period.
Table XI summarizes the portfolio statistics. The sub-period breakdown reveals that performance varied across market regimes: net annualized returns reached 19.1% during the 2017-2019 bull market but compressed to 8.7% during the 2022 inflationary drawdown, when elevated volatility increased turnover and eroded returns through higher transaction costs. This regime sensitivity reinforces the observation that the strategy’s profitability is not uniform and that lower-frequency rebalancing or dynamic cost-awareness could improve net performance in turbulent conditions.
TABLE XI: Portfolio backtest statistics (January 2017 to March 2025).
Statistic Gross Net (10bps)
Annualized Return 22.8% 14.6%
Annualized Sharpe Ratio 1.35 0.84
Maximum Drawdown 12.1% 14.3%
Average Daily Turnover 32%
S&P 500 Buy-and-Hold Return 13.9%
S&P 500 Maximum Drawdown 23.9%
@@ VI-A Interpretation of Results
Within the scope of this evaluation, the integrated model shows consistent improvements across metrics, horizons, and market conditions, with the magnitude of gains varying by component and context in ways that shed light on the mechanisms underlying the model’s predictive performance.
@@
Deeper inspection of the learned graph reveals three structural properties that align with established financial theory. First, edge weights cluster by sector in a manner that goes beyond the initialization prior. Technology stocks (Apple, Microsoft, Salesforce, Netflix) form the tightest cluster with average intra-group weights of 0.81, followed by energy (ExxonMobil, Chevron) at 0.79 and healthcare (Johnson & Johnson, UnitedHealth, Pfizer) at 0.74. Consumer goods and industrials exhibit weaker internal cohesion (0.63 and 0.58 respectively), consistent with the greater heterogeneity of business models within those sectors. Second, the temporal stability of edge weights varies predictably with economic conditions. During the low-volatility period of 2017-2019, the standard deviation of edge weights across consecutive months averaged 0.04, indicating a stable relationship structure. This stability broke down during the March 2020 COVID crash, when the month-over-month standard deviation spiked to 0.11 and cross-sector edge weights increased by an average of 0.18, reflecting the well-documented phenomenon of correlation convergence during systemic stress. By mid-2021, the weight distribution had largely reverted to pre-crisis levels. Third, certain cross-sector edges exhibited persistently high weights that reflect supply-chain or macroeconomic linkages rather than sector proximity. The learned weight between ExxonMobil and Caterpillar (0.67, the highest cross-sector pair) likely captures the shared sensitivity of energy and heavy-industry firms to commodity cycles and capital expenditure trends. Table XII lists the five highest and five lowest learned edge weights to illustrate the range of discovered relationships.
TABLE XII: Top-5 and bottom-5 learned edge weights after training.
Stock $i$ Stock $j$ Weight Same Sector
Apple Microsoft 0.86 Yes
ExxonMobil Chevron 0.82 Yes
Microsoft Salesforce 0.78 Yes
J&J UnitedHealth 0.76 Yes
ExxonMobil Caterpillar 0.67 No
Netflix Chevron 0.19 No
Pfizer Nike 0.21 No
Verizon Salesforce 0.23 No
Boeing Coca-Cola 0.24 No
McDonald’s Netflix 0.26 No
The top weights are dominated by same-sector pairs whose business fundamentals are closely linked, while the bottom weights connect companies with minimal overlap in revenue drivers, customer bases, or macroeconomic exposures. The one exception among the top five, ExxonMobil-Caterpillar, reflects a learned cross-sector dependency that the initialization prior alone would not have produced. These patterns suggest that the edge refinement mechanism successfully discovers economically meaningful relationships that go beyond the sector and correlation priors used for initialization.
Sentiment integration appears particularly valuable around information events. The 25% improvement during earnings seasons suggests that social media sentiment may carry early signals not yet fully reflected in contemporaneous prices. Sector-level heterogeneity aligns with expectations: consumer-facing companies show stronger sentiment-price relationships (14% MAPE reduction) compared to utilities (6%), likely reflecting differential retail investor engagement.
@@ VI-B Implications
For practitioners, the model’s directional accuracy could support more informed trading decisions, particularly for short-term strategies where predicting price direction is often more important than predicting magnitude. The ability to maintain lower error during volatile conditions could assist risk management, as these are precisely the periods when forecast reliability matters most for position sizing and drawdown control. The graph structure provides interpretable inter-stock relationships that may inform portfolio diversification, as the learned edge weights can surface dependencies that traditional correlation matrices might obscure. These practical implications remain conditional on the caveats discussed in Section VI-C, including survivorship bias in the evaluation universe and simplified transaction cost assumptions.
From a theoretical perspective, performance gains from sentiment integration suggest that markets may not fully incorporate qualitative information immediately, a finding consistent with bounded efficiency [1]. The 25% improvement during earnings seasons suggests that social media may contain early signals about investor reactions that have not yet been absorbed into price. The effectiveness of graph-based modeling is consistent with the view that markets function as complex, interconnected systems rather than collections of independent securities, though the limited evaluation universe prevents strong generalization of this observation.
@@ VI-C Limitations
The most consequential concern is survivorship bias in stock selection. The 20 stocks were chosen based on current S&P 500 membership and data availability, meaning that companies which failed, were acquired, or were delisted between 1982 and 2025 are absent from the dataset. Because the selected universe consists exclusively of companies that ultimately succeeded, the dataset omits firms that experienced bankruptcy, prolonged decline, or delisting, which are typically less predictable and more volatile. Reported performance is therefore likely inflated relative to what a real-time investor would have experienced using historical index constituents, and the results should be interpreted as demonstrating the architecture’s effectiveness on a stable set of well-established, liquid equities rather than as a guarantee of generalizable performance across the full market. Constructing a point-in-time universe from historical S&P 500 membership records, including reconstitution events, would provide a more realistic evaluation and is identified as a priority for future work. Relatedly, the 20-node graph structure, while demonstrating meaningful contribution in ablation studies (15% improvement), represents a limited cross-section of the market. Larger universes might better capture complex market topology, and whether the learned edge weights retain predictive value as the graph grows from 20 nodes to hundreds remains an open question. The current design prioritizes temporal depth over cross-sectional breadth, serving as a controlled environment to validate the graph-based modeling approach before scaling to broader equity populations.
Beyond stock selection, data-related assumptions also affect generalizability. Edge weight initialization relies on correlations computed over the training period (1982-2010), but market correlations are inherently non-stationary. Relationships that held during this window may weaken or reverse by the test period. The learnable edge refinement mechanism partially compensates by adapting weights during training, though fundamental correlation regime shifts could still limit generalization. Similarly, X (formerly Twitter) data availability from 2007 onward means sentiment features are set to zero for the 1982-2006 training period, creating a pronounced imbalance: the sentiment fusion mechanism trains on only four years of multimodal data (2007-2010) while the price-based components benefit from 28 years of signal. Although the adaptive gating mechanism mitigates this by learning to suppress the sentiment channel when its input is zero, the limited multimodal training window may constrain the model’s ability to fully exploit sentiment-price interactions. The ablation study (Section V-E) confirms that the architecture retains strong performance without sentiment (MAPE 0.88%), and the reported 10% improvement from sentiment integration should be understood as a lower bound that could increase with a longer sentiment training history. Sentiment analysis also relies on English-language content from a single social media platform, and changes to API access policies may affect reproducibility.
@@ VI-D Future Directions
Multiple directions merit investigation. A critical next step is evaluating the framework on a point-in-time stock universe reconstructed from historical index membership records, which would eliminate survivorship bias and provide a more realistic assessment of real-world performance. Beyond bias correction, expanding the stock universe to the full S&P 500 or international markets would test whether the graph-based architecture scales effectively and whether inter-stock dependencies remain informative as the number of nodes grows. Incorporating multilingual sentiment sources and alternative platforms (Reddit, financial news, earnings call transcripts) would broaden the information base and reduce single-source dependency. On the architectural side, developing more efficient attention mechanisms, such as sparse or linear attention variants, could enable real-time deployment for higher-frequency trading applications. Regime-shift detection mechanisms that adjust model behavior during structural market transitions (e.g., policy changes or liquidity crises) represent a natural extension of the gating mechanism already present in the model. Finally, integrating the forecasting framework with portfolio optimization and formal risk management objectives could bridge the gap between prediction accuracy and realized economic value.
@@ VII Conclusion
Taken together, these findings provide preliminary evidence that jointly modeling inter-stock dependencies and investor sentiment can improve forecast accuracy relative to architectures that address either dimension in isolation. Several limitations temper the strength of this conclusion: the evaluation covers 20 large-cap U.S. equities selected with survivorship bias, the sentiment training window spans only a fraction of the full training period, and the economic backtest applies simplified transaction cost assumptions. Whether these improvements extend to broader, historically reconstructed equity universes and withstand realistic trading frictions remains an open question that the current study cannot resolve. Future work should prioritize evaluation on a point-in-time stock universe, incorporate multilingual and multi-platform sentiment sources, investigate efficient attention variants for real-time deployment, and integrate the forecasting pipeline with portfolio optimization and formal risk management objectives.